OpenStudy (anonymous):

(i really need help with this!will give medals) solve the system by graphing: {-3x-y=-10 {4x-4y=8 my question here is how i get the coords for the 2 lines that intersect from the system above. i know the answer is the point where the two lines intersect, but how do i get the lines? please and thankyou in advance for help!

OpenStudy (abb0t):

Solve by graphing? Well, what you CAN do is rearrange them in the form of a line: \(\sf \color{red}{y=mx+b} \)

OpenStudy (abb0t):

graph them like that and see where the points interect.

OpenStudy (anonymous):

alright thanks!

OpenStudy (abb0t):

Or you can also use substitution method or elimination method to find the points.

OpenStudy (anonymous):

you make this sound so much easier than it looks at the moment. haha, right now i just rearranged them into y=mx+b but im still confused as to pulling a line out of them. how would i use the substitution method for this? ive always been better at that

OpenStudy (abb0t):

Well, you know that y=mx+b is a line, yes? And if you rearrange the first one, you should get y = -3x+10

OpenStudy (abb0t):



OpenStudy (anonymous):

okay so this is starting to make since now. but the -3x, what is that used for when directing the line? m is slope right?

OpenStudy (abb0t):

That is your first line. Now, rearrange the second one y=x +2

OpenStudy (abb0t):

Yes, m = slope = \(\frac{rise}{run}\)

OpenStudy (anonymous):

alright then if you were to put the second one up would it have a slope of 1 and intersect the y axis at +2?

OpenStudy (abb0t):



OpenStudy (anonymous):

awesome! thanks so much, i'm pretty sure i get it now